Write a rational function f(x) with asymptotes at y=7,x=-4 and a hole at x=2.

So far i got f(x)= (7x^2-14)/)(x+2)(x-2))

Answer by richard1234(7193)

You should check x = -4 and x = 2. If x = -4, then f(-4) = 98/12 which is not undefined, so there is no asymptote there. At x = 2, f(2) = 14/0, which means there is an asymptote, not a hole. Basically, check Wolfram Alpha or some graphing utility first.

You are right that the ratio of the leading coefficients is 7 and that the numerator/denominator have the same degree. For x = -4, we want to obtain f(-4) = k/0 where k is non-zero. For x = 2, we want f(2) = 0/0 but to exist.
Start by putting x+4 and x-2 in the denominator. Let where has degree at most 1. Then and . Try and you will see both conditions are satisfied. Hence, an example function that satisfies all of the constraints is